OpenStudy (anonymous):

Decimals how do you write a decimal that keeps going forever but doesn't repeat? How do you write a number that goes on forever but does repeat(has a pattern?) how do you turn both of those into a fraction and percent? How do you round them (in finance or other?) how many ways can you round? And how exactly do you write them on paper? I've seen at least a few different ways that people round them and I don't understand.

OpenStudy (anonymous):

Please be as detailed as possible

OpenStudy (zpupster):

how do you write a decimal that keeps going forever but doesn't repeat? \[\pi = 3.1428571428571.....\] or 22/7 is a number that goes on forever and does not repeat. how do you write a decimal that keeps going forever but doesn't repeat?

OpenStudy (anonymous):

I'm asking how you would write it on paper. How would you put it? How would you round it? Everything I asked above

OpenStudy (zpupster):

\[1/3=0.\overline{33333333} \ or \\ 1/7= \overline{142857}\]

OpenStudy (zpupster):

how many ways can you round? here are a couple of links to explain: https://www.mathsisfun.com/rounding-numbers.html https://www.mathsisfun.com/numbers/rounding-methods.html

OpenStudy (zpupster):

how do you turn both of those into a fraction and percent? 1/2 divide numerator --number on top by denominator-- number on bottom = 0.5 to get a per cent multiply by 100 =50%

OpenStudy (zpupster):

How do you round them (in finance or other? http://www.eecis.udel.edu/~breech/contest.inet.fall.07/problems/bankers-rounding.html

OpenStudy (phi):

how do you write a decimal that keeps going forever but doesn't repeat? you obviously can't write all the decimals because you will never finish. but you can give a name to it and what it represents. for example: the ratio of any circle's circumference divided by its diameter that is a number, and it turns out to be a number with decimals that go on forever. or you can write such a number by saying it is the solution to some specific equation.

OpenStudy (anonymous):

Thank you both for your help!